OpenStudy (anonymous):

At his job, Rob works 6 hours on Monday, 7 hours on Tuesday, and 2 hours on Friday. He receives a $15 bonus, which causes his weekly paycheck to total $153.75. What is Rob's hourly pay?

OpenStudy (anonymous):

6+7+2=15 153.75-15=138.75 138.75/15=9.25/hour

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np